OpenStudy (anonymous):

Which is the equation of the line with slope 3 that contains point (-1, 5)? y – 5 = 3(x + 1) y – 1 = 3(x + 5) y – 1 = 3(x – 5) y – 5 = 3(x - 1)

OpenStudy (anonymous):

x=5, y=-1, m=3 \[y-y _{1}=m(x-x _{1})\] y-5 = 3(x-(-1)) y-5 = 3(x+1)